# Anomalous Diffusion in Driven Electrolytes due to Hydrodynamic Fluctuations

### v1

@@ Abstract
Crossover Time-scales.—It is instructive to examine the behaviour of crossover time-scales for the possible scenarios. The crossover from the ballistic regime to the first anomalous regime occurs at a time-scale $\tau^{(d)}_{\rm ba}$ that can be obtain via $\left\langle{{\bm{v}}^{2}}\right\rangle(\tau^{(d)}_{\rm ba})^{2}={\cal D}_{z_{1}}^{3-d/2}(\tau^{(d)}_{\rm ba})^{3-d/2}$, where we will use input from Eqs. (10) and (14). While for $d=2$, this does not yield a result, we find $\tau^{(d)}_{\rm ba}=a^{2}/D$ for $d>2$ and $\tau^{(d)}_{\rm ba}=L^{2}/D$ for $d<2$. This implies that the ballistic regime only exits for $d>2$, as the velocity fluctuations diverge with the system size in $d<2$.
@@
The crossover time from ballistic regime to diffusive regime, $\tau^{(d)}_{\rm bd}$, can be obtained via $\left\langle{{\bm{v}}^{2}}\right\rangle(\tau^{(d)}_{\rm bd})^{2}={\cal D}_{\rm sc}\tau^{(d)}_{\rm bd}$. Note that this crossover exists only for $d\geq 4$. Using Eqs. (10), (18), and (20), the calculation yields $\tau^{(d)}_{\textrm{b}\textrm{d}}=C_{0}^{1/2}\lambda_{\rm e}^{-1}a^{d/2}$ for $d>4$ and $\tau^{(4)}_{\textrm{b}\textrm{d}}=C_{0}^{1/2}\lambda_{\rm e}^{-1}a^{2}\sqrt{\ln(L/a)}$ for $d=4$. For $d<4$, a long-time crossover occurs at time $\tau_{L}^{(d)}$ when the MSD in the second anomalous regime is saturated by the system size $L$. Using $L^{2}={\cal D}_{z_{2}}^{4/d}(\tau_{L}^{(d)})^{4/d}$ and Eq. (15), the calculation gives $\tau_{L}^{(d)}=C_{0}^{1/2}\lambda_{\rm e}^{-1}L^{d/2}$ for $d<4$. Note that a similar calculation by using the divergent self-consistent effective diffusion equation from Eq. (19) yields the same result, namely, $\tau_{L}^{(d)}=L^{2}/{\cal D}_{\rm sc}=C_{0}^{1/2}\lambda_{\rm e}^{-1}L^{d/2}$ for $d<4$. Finally, since in $d=2$, the anomalous diffusion exponents for both regimes are the same as the ballistic dynamics exponent, i.e. $\alpha_{1}=\alpha_{2}=2$, the ballistic regime is expected to govern all time scales in this case (see Fig. 2b).
@@
Concluding Remarks.—The stochastic fluctuations of tracers in a bulk electrolyte have been found to exhibit a wide range of dynamical regimes and crossover scales that depend strongly on dimensionality. The existence of two anomalous regimes is particularly interesting, with the anomalous diffusion exponent of the first regime $\alpha_{1}=3-d/2$ being related to the anomalous dynamics of catalytically active colloids [32, 26]. Anomalous dynamics can also arise in anisotropic active suspensions [33]. Due to the important role that ionic currents play in molecular sensing technologies [3, 4], it will be important to investigate how such strong and anomalous fluctuations can influence these sensing devices. A starting point can be to incorporate hydrodynamic and ionic effects in simple models of molecular sensing that use polymer translocation [34].
@@ Acknowledgements.
I acknowledge discussions with Sriram Ramaswamy, and support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.

### v2

@@ Abstract
Crossover time-scales.—It is instructive to examine the behaviour of crossover time-scales for the possible scenarios. The crossover from the ballistic regime to the first anomalous regime occurs at a time-scale $\tau^{(d)}_{\rm ba}$, which can be obtained via $\left\langle{{\bm{v}}^{2}}\right\rangle(\tau^{(d)}_{\rm ba})^{2}={\cal D}_{z_{1}}^{3-d/2}(\tau^{(d)}_{\rm ba})^{3-d/2}$, where we will use input from Eqs. (11) and (15). While for $d=2$, this does not yield a result, we find $\tau^{(d)}_{\rm ba}=a^{2}/D$ for $d>2$ and $\tau^{(d)}_{\rm ba}=L^{2}/D$ for $d<2$. This implies that the ballistic regime only exits for $d>2$, as the velocity fluctuations diverge with the system size in $d<2$ (see above).
@@
The crossover time from ballistic regime to diffusive regime, $\tau^{(d)}_{\rm bd}$, can be obtained via $\left\langle{{\bm{v}}^{2}}\right\rangle(\tau^{(d)}_{\rm bd})^{2}={\cal D}_{\rm sc}\tau^{(d)}_{\rm bd}$. Note that this crossover exists only for $d\geq 4$. Using Eqs. (11), (19), and (21), the calculation yields $\tau^{(d)}_{\textrm{b}\textrm{d}}=C_{0}^{1/2}\lambda_{\rm e}^{-1}a^{d/2}$ for $d>4$ and $\tau^{(4)}_{\textrm{b}\textrm{d}}=C_{0}^{1/2}\lambda_{\rm e}^{-1}a^{2}\sqrt{\ln(L/a)}$ for $d=4$. For $d<4$, a long-time crossover occurs at time $\tau_{L}^{(d)}$ when the MSD in the second anomalous regime is saturated by the system size $L$. Using $L^{2}={\cal D}_{z_{2}}^{4/d}(\tau_{L}^{(d)})^{4/d}$ and Eq. (16), the calculation gives $\tau_{L}^{(d)}=C_{0}^{1/2}\lambda_{\rm e}^{-1}L^{d/2}$ for $d<4$. Note that a similar calculation using the divergent self-consistent effective diffusion coefficient from Eq. (20) yields the same result, namely, $\tau_{L}^{(d)}=L^{2}/{\cal D}_{\rm sc}=C_{0}^{1/2}\lambda_{\rm e}^{-1}L^{d/2}$ for $d<4$. Finally, since in $d=2$, the anomalous diffusion exponents for both regimes are the same as the ballistic dynamics exponent, i.e. $\alpha_{1}=\alpha_{2}=2$, the ballistic regime is expected to govern all time scales in this case (see Fig. 2b). Note that the adiabatic elimination of the charge density due to Debye screening relies on the Debye relaxation time $1/(D\kappa^{2})$ being much shorter than the other relevant timescales such as $1/\lambda_{\rm e}$ or the various crossover scales. This is guaranteed by the strong electrolyte condition, as can be seen for example by noting that for physiologically relevant salt concentrations $\kappa^{-1}\sim 1$ nm giving rise to a Debye relaxation time $\sim 1$ ns.
@@
Concluding remarks.—The stochastic fluctuations of tracers in a bulk electrolyte have been found to exhibit a wide range of dynamical regimes and crossover scales that depend strongly on dimensionality. The existence of two anomalous regimes is particularly interesting, as they correspond to a crossover between a regime where density fluctuations directly result in anomalous dynamics (characterized by $\alpha_{1}=3-d/2$) to a regime where advection by the fluid flow changes the nature of the anomalous dynamics (characterized by $\alpha_{2}=4/d$) albeit driven by the same density fluctuations. This phenomenon is intrinsically related to the anomalous dynamics of catalytically active colloids where only the first regime was originally considered [Golestanian2009b, Golestanian2019phoretic], due to similarities in the interplay between density fluctuations and dynamics under non-equilibrium driving forces and fluxes. Anomalous dynamics can also arise in anisotropic active suspensions [sriram0]. Because of the important role that ionic currents play in molecular sensing technologies [Ying2022, Ritmejeris2024], it will be important to investigate how such strong and anomalous fluctuations can influence these sensing devices. A starting point can be to incorporate hydrodynamic and ionic effects in simple models of molecular sensing that use polymer translocation [Cohen2012].
@@ Acknowledgements.
The author acknowledges discussions with Sriram Ramaswamy and support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.